# Optimal e-value testing for properly constrained hypotheses

### v1

@@ Appendix D Proof of Proposition 1
In $\mathbb{R}^{m}$, a straight line $r$ intersecting the interior of a convex set $C$ is fully contained in $C$ if, and only if, it is parallel to all the supporting hyperplanes of the closure of $C$.
@@
First, note that since $r$ intersects the interior of $C$, it intersects the supporting hyperplanes of the closure of $C$ in at most one point, and hence cannot lie on one of these hyperplanes. Now, if it intersects one of these hyperplanes in only one point, then it cannot be fully contained in $C$. On the other hand, if $r$ is not contained in $C$, $r$ must intersect the boundary of $C$. Since every point of the boundary of $C$ is contained in at least one supporting hyperplane of the closure of $C$, we conclude that $r$ must intersect this hyperplane. ∎
@@
Let $C\subseteq\mathbb{R}^{m}$ be a convex set with $0\in\operatorname*{int}C$. Let $V=\{v\in\mathbb{R}^{m}\,:\,\operatorname*{lin}\{v\}\subseteq C\}$. Then, $V$ is a subspace of $\mathbb{R}^{m}$. Moreover, there is a (possibly non orthogonal) basis $\mathcal{B}$ of $V^{\perp}$ such that $\inf_{w\in C}v\cdot w>-\infty$, for all $v\in\mathcal{B}$.
@@
Clearly, $0\in V$. Now, let $v$ and $v^{\prime}$ belong to $V$. Since $C$ is convex, the fact that both $\operatorname*{lin}\{v\}$ and $\operatorname*{lin}\{v^{\prime}\}$ lie in $C$ implies that $\operatorname*{lin}\{v,v^{\prime}\}\subseteq C$. In particular, for any real coefficients $\alpha$ and $\alpha^{\prime}$, we have $\alpha v+\alpha^{\prime}v^{\prime}\in V$. Hence, $V$ is a subspace of $\mathbb{R}^{m}$.
Let $S$ denote the set of all the normal vectors to the supporting hyperplanes of the closure of $C$, and let $U=\operatorname*{lin}S$. Since $0\in\operatorname*{int}C$, by Lemma 27, $U=V^{\perp}$. By definition of $U$, it is hence possible to find a basis of $V^{\perp}$ made of elements in $S$. Now, let $v\in S$. Then there is a supporting hyperplane $\pi$ of the closure of $C$ that is orthogonal to $v$. This means that there is $\alpha\in\mathbb{R}$ such that $\pi=\{w\in\mathbb{R}^{m}\,:\,v\cdot w=\alpha\}$. Since the origin is in the interior of $C$, it must be that $\alpha\neq 0$. If $\alpha<0$, then $C\subseteq\{w\in\mathbb{R}^{m}\,:\,v\cdot w\geq\alpha\}$, by definition of supporting hyperplane and since $0\in C$. In such a case it is clear that $\inf_{w\in C}v\cdot w\geq\alpha>-\infty$. On the other hand, if $\alpha>0$, we replace $v$ with $-v$ in the basis, and we go back to the case $\alpha<0$. ∎
Lemma 29.
Let $\mathcal{Y}\subseteq\mathbb{R}^{m}$ be any set. Let $C$ denote the convex hull of $\mathcal{Y}$. Assume that for some $\varepsilon>0$, the closed 1-ball of radius $\varepsilon$ centred in 00 is included in $C$. Define $L:\mathbb{R}^{m}\to(0,1]$ as
$L(y)=\frac{\varepsilon}{\varepsilon+\|y\|_{1}}\,.$
Let $V=\{v\in\mathbb{R}^{m}\,:\,\operatorname*{lin}\{v\}\subseteq C\}$ and $\Pi$ be the orthogonal projection from $\mathbb{R}^{m}$ to $V^{\perp}$. For any $y\in\mathbb{R}^{m}$, let $\mathcal{P}_{y}^{\circ}$ denote the set of measures in $\mathcal{P}_{\mathcal{Y}\cup\{y\}}$ whose support has finitely many elements, and let $\mathcal{H}_{y}^{\circ}=\{P\in\mathcal{P}_{y}^{\circ}\,:\,\langle P,{Y}\rangle=0\}$. Then, $\sup_{P\in\mathcal{H}_{y}^{\circ}}P(\{y\})\geq L(\Pi(y))$, for any $y\in\mathbb{R}^{m}$.
@@
For $y\in\mathbb{R}^{m}$, let $S_{y}=\sup_{P\in\mathcal{H}_{y}^{\circ}}P(\{y\})$. We will proceed in two steps.
  1. 1.
    First, we show that $S_{y}\geq L(y)$, for any $y\in\mathbb{R}^{m}$.
  2. 2.
    Then, we prove the tighter bound $S_{y}\geq L(\Pi(y))$, for any $y\in\mathbb{R}^{m}$.
For $y=0$, these claims are trivially true (since $\delta_{0}\in\mathcal{H}_{0}^{\circ}$). We will henceforth assume that $y\neq 0$.
  1. 1.
    Fix $y\in\mathbb{R}^{m}$ non-zero. Since the centred closed ball of radius $\varepsilon$ is in $C$, $z=-\varepsilon y/\|y\|_{1}\in C$. Let $\mathcal{P}_{y}^{\circ}$ denote the set of probability measures on $\mathbb{R}^{m}$ whose support has finitely many elements and is included in $\mathcal{Y}$. By Lemma 22, there is $P\in\mathcal{P}_{y}^{\circ}$, such that $\langle P,Y\rangle=z$. Now consider $P^{\prime}=\frac{\varepsilon}{\varepsilon+\|y\|_{1}}\delta_{y}+\frac{\|y\|_{1}}{\varepsilon+\|y\|_{1}}P$. Clearly, $P^{\prime}\in\mathcal{P}_{y}^{\circ}$. Moreover, it is easily checked that $\langle P^{\prime},Y\rangle=0$, so $P^{\prime}\in\mathcal{H}_{y}^{\circ}$. Since $P^{\prime}(\{y\})\geq L(y)$, it follows that $S_{y}\geq L(y)$.
  2. 2.
    Fix a non-zero $y$ in $\mathbb{R}^{m}$. By Lemma 28, $V$ is a subspace. Let $y_{\perp}=\Pi(y)$ and $y_{V}=y-y_{\perp}$. Then, $y_{V}\in V$. Thus, for any $u>0$ we have that $z_{u}=-uy_{V}\in C$. By Lemma 22, there is $P_{u}\in\mathcal{P}_{y}^{\circ}$ such that $\langle P_{u},Y\rangle=z_{u}$. Let $y_{u}=\frac{u}{1+u}y_{\perp}$ and $P^{\prime}_{u}=\frac{u}{1+u}\delta_{y}+\frac{1}{1+u}P_{u}\in\mathcal{P}_{y}^{\circ}$. Clearly, $\langle P_{u}^{\prime},Y\rangle=\frac{u}{1+u}(y-y_{V})=y_{u}$. Fix an arbitrary $\alpha>0$. From what we have already shown, we know that there is $P^{\prime\prime}_{u}\in\mathcal{H}_{y_{u}}^{\circ}$ such that $P^{\prime\prime}_{u}(\{y_{u}\})>L(y_{u})-\alpha$. Denoting $p=P^{\prime\prime}_{u}(\{y_{u}\})$, let $P^{\prime\prime}_{u}=p\delta_{y_{u}}+(1-p)P^{\prime}$, with $P^{\prime}\in\mathcal{P}_{y}^{\circ}$. Let $\hat{P}_{u}=pP^{\prime}_{u}+(1-p)P^{\prime}$. Then, $\hat{P}_{u}\in\mathcal{P}_{y}^{\circ}$ and $\langle\hat{P}_{u},Y\rangle=py_{u}+(1-p)\langle P^{\prime},Y\rangle=\langle P^{\prime\prime}_{u},Y\rangle=0$, since $P^{\prime\prime}_{u}\in\mathcal{H}_{y_{u}}^{\circ}$. So, $\hat{P}_{\mu}\in\mathcal{H}_{y}^{\circ}$ and $\hat{P}_{u}(\{y\})\geq pP^{\prime}_{u}(\{y\})\geq\frac{up}{1+u}>\frac{u}{1+u}(L(y_{u})-\alpha)$. The whole construction is valid for any $u>0$ and $\alpha>0$. So, $S_{y}\geq\lim_{\alpha\to 0}\lim_{u\to+\infty}\frac{u}{1+u}(L(y_{u})-\alpha)=L(\Pi(y))$, since $L$ is continuous and $y_{u}\to y_{\perp}=\Pi(y)$ for $u\to+\infty$.
This concludes the proof. ∎
Lemma 30.
Let $\mathcal{Y}\subseteq\mathbb{R}^{m}$ be any set. Let $C$ denote the convex hull of $\mathcal{Y}$, and assume that $0\in\operatorname*{int}C$. Let $\mathcal{P}_{1}=\{P\in\mathcal{P}_{\mathcal{Y}}\,:\,\langle P,\|Y\|_{1}\rangle\leq 1\}$, and =$\mathcal{H}=\{P\in\mathcal{P}_{1}\,:\,\langle P,Y\rangle=0\}$. Let $\mathcal{H}^{\circ}$ denote the set of measures in $\mathcal{H}$ whose support has finite cardinality. Then, there is a continuous function $A:\mathcal{Y}\to[1,+\infty)$, such that, for any $y\in\mathcal{Y}$, $A(y)\sup_{P\in\mathcal{H}^{\circ}}P(\{y\})\geq 1$, and $\langle P,A\rangle<+\infty$ for all $P\in\mathcal{H}$.
Proof.
Since $0\in\operatorname*{int}C$, we can fix $\varepsilon>0$ such that the centred closed 1-ball of radius $\varepsilon$ is in $C$. Let $L$, $V$, and $\Pi$ be as in Lemma 29. For $y\in\mathcal{Y}$, let $A(y)=1/L(\Pi(y))=1+\|\Pi(y)\|_{1}/\varepsilon$. Then, $A:\mathcal{Y}\to[1,+\infty)$ is continuous and, for any $y\in\mathcal{Y}$, $A(y)\sup_{P\in\mathcal{H}^{\circ}}P(\{y\})\geq 1$ by Lemma 29.
Let $p\leq m$ be the dimension of $V^{\perp}$. By Lemma 28, we can fix a basis =$\mathcal{B}=\{v_{1},\dots,v_{m}\}$ of $\mathbb{R}^{m}$, such that $\{v_{p+1},\dots v_{m}\}$ is a basis of $V$ and $\{v_{1},\dots,v_{p}\}$ is a basis of $V^{\perp}$, where for each $i\leq p$ we have that $\inf_{w\in C}w\cdot v_{i}>-\infty$. For $y\in\mathbb{R}^{m}$, let $\|y\|_{\mathcal{B}}=\sum_{i=1}^{m}|v_{i}\cdot y|$. Since $\mathcal{B}$ is a basis, this is a norm on $\mathbb{R}^{m}$. As all norms on finite dimensional spaces are equivalent, we can find a constant $K>0$ such that, for any $y\in\mathbb{R}^{m}$, $\|y\|_{1}\leq K\|y\|_{\mathcal{B}}$. It follows that, for any $y\in\mathcal{Y}$,
$A(y)\leq 1+\frac{K}{\varepsilon}\|\Pi(y)\|_{\mathcal{B}}=1+\frac{K}{\varepsilon}\sum_{i=1}^{p}|v_{i}\cdot y|\,.$
Now, for any $i\leq p$ and any $P\in\mathcal{H}$, we have that $\langle P,v_{i}\cdot Y\rangle=0$. Since $v_{i}\cdot Y$ is bounded from below (as $Y$ is supported in $C$), it follows that $\langle P,|v_{i}\cdot Y|\rangle$ is finite. Hence, $\langle P,A\rangle<+\infty$. ∎
Proposition 1.
Let $\mathcal{H}$ be a properly constrained hypothesis on $\mathcal{X}$. There is a continuous function $A:\mathcal{X}\to[1,+\infty)$ such that $\langle P,A\rangle<+\infty$ for every $P\in\mathcal{H}$, $\sup_{P\in\mathcal{H}}P(\{x\})\geq 1/A(x)$ for all $x\in\mathcal{X}$, and $E\preceq A$ for any $E\in\mathcal{E}_{\mathcal{H}}$.
Proof.
By Lemma 5, we can find a minimal constraint $\Phi:\mathcal{X}\to\mathbb{R}^{m}$ for $\mathcal{H}$. Let $\mathcal{Y}=\Phi(\mathcal{X})$ and $C=\operatorname*{conv}\mathcal{Y}$. Define =$\mathcal{H}^{\prime}=\{\Phi^{\#}P\,:\,P\in\mathcal{H}\}$, where $\Phi^{\#}P$ is the push-forward of $P$ under $\Phi$. Let $\mathcal{H}^{\circ}$ be the set of measures in $\mathcal{P}_{\mathcal{Y}}$ with zero mean and finite support, which is non-empty by Lemma 22. Fix any $P^{\circ}\in\mathcal{H}^{\circ}$. For each $y\in\operatorname*{Supp}P^{\circ}$, $\xi_{y}=\Phi^{-1}(\{y\})$ is Borel measurable. Since all these $\xi_{y}$ are disjoint, we can define a Borel measure $\tilde{P}^{\circ}$ on $\mathcal{X}$, supported on $\cup_{y\in\operatorname*{Supp}P^{\circ}}\xi_{y}$, such that $\tilde{P}^{\circ}(\xi_{y})=P^{\circ}(\{y\})$. It’s easy to verify that $P^{\circ}=\Phi^{\#}\tilde{P}^{\circ}$ and $\tilde{P}^{\circ}\in\mathcal{P}_{\Phi}$. So, $\langle\tilde{P}^{\circ},\Phi\rangle=\langle P^{\circ},Y\rangle=0$. Hence, $\tilde{P}^{\circ}\in\mathcal{H}$ and $P^{\circ}\in\mathcal{H}^{\prime}$. We conclude that $\mathcal{H}^{\circ}\subseteq\mathcal{H}^{\prime}$.
Since $\Phi$ is a minimal constraint, $0\in\operatorname*{int}C$. Since for every $P^{\prime}\in\mathcal{H}^{\prime}$ we have that $\langle P^{\prime},Y\rangle=0$ and $\operatorname*{Supp}P^{\prime}\subseteq\mathcal{Y}$, by Lemma 30 there is a continuous $\tilde{A}:\mathcal{Y}\to[1,+\infty)$, such that $\langle P^{\prime},\tilde{A}\rangle<+\infty$ for all $P^{\prime}\in\mathcal{H}^{\prime}$, and $\sup_{P^{\prime}\in\mathcal{H}^{\prime}}P^{\prime}(\{y\})\tilde{A}(y)\geq\sup_{P^{\prime}\in\mathcal{H}^{\circ}}P^{\prime}(\{y\})\tilde{A}(y)\geq 1$, for any $y\in\mathcal{Y}$. Now, let $A=\tilde{A}\circ\Phi$. Clearly, $A$ is a continuous function from $\mathcal{X}$ to $[1,+\infty)$. Moreover, for any $P\in\mathcal{H}$ we have $\langle P,A\rangle=\langle P,\tilde{A}\circ\Phi\rangle=\langle\Phi^{\#}P,\tilde{A}\rangle<+\infty$, since $\Phi^{\#}P\in\mathcal{H}^{\prime}$.
Now, fix $x\in\mathcal{X}$. Let $y=\Phi(x)$ and $\xi=\Phi^{-1}(\{y\})\subseteq\mathcal{X}$. Clearly, $\xi$ is Borel measurable, as $\Phi$ is continuous. Fix any $P^{\prime}\in\mathcal{H}^{\prime}$. By construction, there is $P\in\mathcal{H}$ such that $P^{\prime}=\Phi^{\#}P$. In particular, $P(\xi)=P^{\prime}(\{y\})$. Now, let $P^{\star}$ be the non-negative measure on $\mathcal{X}$ that is equal to $P$ on $\mathcal{X}\setminus\xi$, and gives no mass to $\xi$. Because $P\in\mathcal{H}$, $\int_{\mathcal{X}}\|\Phi\|_{1}\mathrm{d}P^{\star}<+\infty$ and $\int_{\mathcal{X}}\Phi\mathrm{d}P^{\star}=-P(\xi)y$. Let $\hat{P}=P^{\star}+P(\xi)\delta_{x}$. Then, $\hat{P}\in\mathcal{P}_{\Phi}$. Moreover, $\langle\hat{P},\Phi\rangle=P(\xi)y+\langle P^{\star},\Phi\rangle=0$. Hence, $\hat{P}\in\mathcal{H}$, and $\hat{P}(\{x\})=P(\xi)=P^{\prime}(\{y\})$. Thus, for any $P^{\prime}\in\mathcal{H}^{\prime}$, there is $\hat{P}\in\mathcal{H}$ such that $\hat{P}(\{x\})=P^{\prime}(\{y\})$. So, $\sup_{P\in\mathcal{H}}P(\{x\})\geq\sup_{P^{\prime}\in\mathcal{H}^{\prime}}P^{\prime}(\{\Phi(x)\})\geq 1/\tilde{A}(\Phi(x))=1/A(x)$.
Finally, for any e-variable $E$ and any $x\in\mathcal{X}$, we have
$E(x)\leq A(x)\sup_{P\in\mathcal{H}}P(\{x\})E(x)\leq A(x)\sup_{P\in\mathcal{H}}\langle P,E\rangle\leq A(x)\,,$
and so we conclude. ∎

### v3

@@ Appendix C Proof of Proposition 6
Let $\mathcal{H}$ be a properly loosely constrained hypothesis on $\mathcal{X}$ and $\mathcal{X}$ have finite cardinality. Then, for any $Q\in\mathcal{P}_{\mathcal{X}}$, the set $\operatorname*{arg\,max}_{E\in\mathcal{E}_{\mathcal{H}}}\langle Q,\log E\rangle$ is non-empty. Moreover, all its elements coincide on $\operatorname*{Supp}Q$.
@@
Fix $Q\in\mathcal{P}_{\mathcal{X}}$. Since $\mathcal{E}_{\mathcal{H}}$ is compact and $\Phi:E\mapsto\langle Q,\log E\rangle=\sum_{x\in\operatorname*{Supp}Q}Q(\{x\})\log E(x)$ is, upper semi-continuous, concave, and valued on [,+)[$\infty$,$\infty$)[ - ∞ , + ∞ ), we have that $\operatorname*{arg\,max}_{E\in\mathcal{E}_{\mathcal{H}}}\Phi(E)$ is non-empty. Moreover, since the constant function $1$ is an e-variable, the maximum of $\Phi$ is always finite. Now, denoting as $\Pi_{Q}$ the projection of $E\in\mathcal{E}_{\mathcal{H}}$ on $\operatorname*{Supp}Q$ (namely $\Pi_{Q}(E)(x)=E(x)$ if $x\in\operatorname*{Supp}Q$ and 00 otherwise) $\Phi$ is strictly convex on $\Pi_{Q}(\mathcal{E}_{\mathcal{H}})$, which is a convex set. Hence, $\operatorname*{arg\,max}_{E\in\Pi_{Q}(\mathcal{E}_{\mathcal{H}})}\Phi(E)$ is unique, and so all the elements in $\operatorname*{arg\,max}_{E\in\mathcal{E}_{\mathcal{H}}}\Phi(E)$ must coincide on $\operatorname*{Supp}Q$. ∎
Now, we recall a result, which follows directly from Grünwald et al. (2024) and Larsson et al. (2024).
@@
Let $\mathcal{H}$ be a properly loosely constrained hypothesis on $\mathcal{X}$, and assume that $\mathcal{X}$ has finite cardinality. Then, there exists a unique minimiser $P_{Q}\in\mathcal{H}$ of $P\mapsto\mathrm{KL}(Q|P)$, whose support contains $\operatorname*{Supp}Q$. Let $E_{Q}:\mathcal{X}\to\mathbb{R}_{\geq 0}$ be defined as
$E_{Q}(x)=\begin{cases}\tfrac{Q(\{x\})}{P_{Q}(\{x\})}&\text{if $x\in\operatorname*{Supp}Q$;}\\ 0&\text{otherwise.}\end{cases}$
$E_{Q}$ is an e-variable, and $E_{Q}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{\mathcal{H}}}\langle Q,\log E\rangle$. Moreover, if for an e-variable $E$ there is $P\in\mathcal{H}$ such that $E(x)=Q(\{x\})/P(\{x\})$, for all $x\in\operatorname*{Supp}Q$, then $E$ and $E_{Q}$ must coincide on $\operatorname*{Supp}Q$.
We can now show that if $\mathcal{X}$ has finite cardinality and $E$ is a maximal e-variable for a properly loosely constrained $\mathcal{H}$, then $E$ must belong to the dual e-class.
Lemma 28.
Let $\mathcal{X}$ have finite cardinality. Let $\mathcal{H}$ be a properly loosely constrained hypothesis on $\mathcal{X}$, with $\Phi^{\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime}}$ and $\Phi^{\prime\prime}:\mathcal{X}\to\mathbb{R}^{m^{\prime\prime}}$ tight and slack constraints. Let $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}}$ and $\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ be defined as in Proposition 6. If $\hat{E}$ is a maximal e-variable, then $E\in\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$.
@@
First, we notice that following exactly the same steps as in the proof of Lemma 2, we can show that there is $\hat{P}\in\mathcal{H}$ that is fully supported and such that $\langle\hat{P},\hat{E}\rangle=1$. Define $\hat{Q}\in\mathcal{P}_{\mathcal{X}}$ with mass $\hat{Q}(\{x\})=\hat{E}(x)\hat{P}(\{x\})$. By Lemma 27, we have that $\hat{E}\in\operatorname*{arg\,max}_{E\in\mathcal{E}_{\mathcal{H}}}\langle\hat{Q},\log E\rangle$. Moreover, $\operatorname*{Supp}\hat{Q}=\operatorname*{Supp}\hat{E}=\{x\,:\,E(x)>0\}$.
Writing the dual problem for the maximisation of $E\mapsto\langle\hat{Q},\log E\rangle$, a solution must be in the form $E_{\lambda}$ (on $\operatorname*{Supp}\hat{Q}$), with $\lambda\in\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}}$. (This follows from usual dual Lagrangian arguments, and can for instance be obtained following the reasoning in Appendix C.3 of Agrawal et al., 2020). By Lemma 27, it follows that $E_{\lambda}$ and $\hat{E}$ must coincide on $\operatorname*{Supp}\hat{E}$. In particular, $E_{\lambda}\succeq\hat{E}$. Since $\hat{E}$ is maximal, we get $E_{\lambda}=\hat{E}$ on the whole $\mathcal{X}$. ∎
We can now prove Proposition 6.
Proposition 6.
[…]
Then, the optimal e-class for $\mathcal{H}$ exists and coincides with
@@
Let $\Phi:\mathcal{X}\to\mathbb{R}^{m^{\prime}+m^{\prime\prime}}$ be the continuous function whose first $m^{\prime}$ components are $\Phi^{\prime}$ and the last $m^{\prime\prime}$ are $\Phi^{\prime\prime}$. Since $\mathcal{H}$ is loosely properly constrained, $0\in\operatorname*{rel\,int}(\operatorname*{conv}\Phi)$. Moreover, note that $\mathcal{P}_{\Phi}=\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}$, so =$\mathcal{H}^{\prime}=\{P\in\mathcal{P}_{\Phi^{\prime}}\cap\mathcal{P}_{\Phi^{\prime\prime}}\,:\,\langle P,\Phi\rangle=0\}$ is a properly constrained hypothesis, whereof $\Phi$ is a constraint. By construction $\mathcal{H}^{\prime}\subseteq\mathcal{H}$, so $\mathcal{E}_{\mathcal{H}^{\prime}}\supseteq\mathcal{E}_{\mathcal{H}}$. Moreover, $\mathcal{E}_{\mathcal{H}^{\prime}}^{\vee}\supseteq\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$. In particular, since every $E\in\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is a maximal e-variable for $\mathcal{H}^{\prime}$, it must also be a maximal e-variable for $\mathcal{H}$. Hence, by Lemma 1 it is enough to show that $\tilde{\mathcal{E}}_{\mathcal{H}}^{\vee}$ is a majorising e-class.
If $\mathcal{X}$ has finite cardinality, this is Lemma 28. The general case is proved by reproducing step by step the proofs of Proposition 5 and Theorem 1. Indeed, the results that were used for those proofs are essentially based on the compactness of $\Lambda_{\Phi|S}$ (for suitable choices of $S\subseteq\mathcal{X}$). This property holds also for $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}|S}$ (with obvious meaning of notation), as $\tilde{\Lambda}_{\Phi^{\prime},\Phi^{\prime\prime}|S}$ is a closed subset of $\Lambda_{\Phi|S}$ (for the same argument as in Lemma 15). ∎